# Surfel-based Gaussian Inverse Rendering for Fast and Relightable Dynamic Human Reconstruction from Monocular Videos

### v2

@@ 3 Our Method
Figure 2: Overview of our pipeline. We define the radiance ($c_{k}$) and materials ($a_{k},r_{k},m_{k}$) at the canonical space as canonical PBR-aware 2DGS and deform them to the world space via Linear Blend Skinning (LBS). We first optimize with Image reconstruction loss to get the initial shape of the clothed avatar. Based on the rough shape, we optimize the Gaussian attributes with Physically Based Rendering. To model the shadow effect and decouple the materials from lighting, we propose to approximate the calculation of occlusion from the template mesh. Additionally, we apply regularization, i.e., smooth loss, white loss, and normal consistency loss to get a plausible solution for the PBR materials.
In this section, we first introduce the background of surfel-based representation [33, 32]. Then, we describe our framework for relightable dynamic clothed avatar representation and physically-based rendering. The clothed avatar is built upon surfel-based representation [33] with an articulated template [40] based on SMPL [39]. Motivated by the success of 2DGS for high-quality geometry recovery, we propose a PBR-aware 2DGS that incorporates the lighting, roughness, metallics, and albedo into the 2DGS for monocular videos. We also propose a solution for ambient occlusion calculation to enhance the rendering realism and better decoupling of the material and lighting. Finally, we describe our progressive training strategy to achieve geometrically accurate surface reconstruction and accurate albedo estimation. We visualize the overall pipeline of our method in Fig. 2.
@@ 3.1 A revisit of 2DGS
Unlike 3DGS [32], 2DGS [33] adopt flat 2D Gaussians embedded in 3D space for scene representation. The primitive distributes densities within a planar disk (also known as surfel), which makes the primitive better alignment with surfaces. This feature can improve the quality of our geometry reconstruction and achieve better material estimation. In 2DGS [33], each primitive is characterized by a central points $\mathbf{p}_{k}\in\mathbb{R}^{3}$ , two principal tangential vectors $\mathbf{t}_{u},\mathbf{t}_{v}\in\mathbb{R}^{3}$ and two scaling factors $s_{u},s_{v}\in\mathbb{R}$. The scaling factors correspond to the variances of the 2D Gaussian.
@@
$\mathbf{P}(u,v)=\mathbf{p}_{k}+s_{u}\mathbf{t}_{u}u+s_{v}\mathbf{t}_{v}v=[\mathbf{R}_{k}\mathbf{S}_{k},\mathbf{p}_{k}](u,v,1,1)^{T}$ (1)
where $\mathbf{R}_{k}=[\mathbf{t}_{u},\mathbf{t}_{v},\mathbf{t}_{u}\times\mathbf{t}_{v}]\in SO(3)$ denotes the rotation of each surfel, $\mathbf{t}_{u}\times\mathbf{t}_{v}$ denotes the normal orientation of each surfel, and the diagonal matrix $\mathbf{S}_{k}=\text{diag}(s_{u},s_{v},0)\in\mathbb{R}^{3\times 3}$ denotes the scale of each surfel. Each 2D Gaussian primitive also carries an opacity $o$ and a view-dependent appearance $\mathbf{c}$.
For each point $\mathbf{u}=(u,v)$ in the local coordinate, its 2D Gaussian value can be calculated by
$\mathcal{G}(\mathbf{u})=\exp(-\frac{u^{2}+v^{2}}{2})$ (2)
Then, 2D Gaussians are sorted based on the depth of their center and organized into tiles. To render a pixel $x$, we then cast a ray into the scene to compute the intersection $\mathbf{u}(x)$ on the splats and accumulate the texture function with volumetric alpha blending:
$\mathbf{c}(x)=\sum_{i=1}\mathbf{c}_{i}o_{i}\mathcal{G}_{i}(\mathbf{u}(x))\prod_{j=1}^{i-1}(1-o_{j}\mathcal{G}_{j}(\mathbf{u}(x)))$ (3)
@@ 3.2 Clothed humans avatars as dynamic surfels animated by template models
We follow the recent advances [40, 100, 44, 41] to model the human as animatable Gaussians. Specifically, we define the PBR-aware 2DGS in a canonical space and transform them into world space via linear blend skinning (LBS). The Gaussians are animated by a human template model [39] to achieve avatar animation. In contrast to them, we choose the surfel-based primitive [33] as our representation for better geometry accuracy. Furthermore, we define additional parameters for intrinsic properties and lighting on each 2D Gaussian to enable physically based rendering.
@@
$\mathbf{G}=\{\mathbf{p}_{k},\mathbf{R}_{k},\mathbf{S}_{k},\mathbf{c}_{k},o_{k},\mathbf{a}_{k},r_{k},m_{k}\}_{k=1}^{N_{\text{gs}}}$ (4)
@@
$N_{\text{gs}}$ is the number of the gaussians. $\mathbf{p}_{k},\mathbf{R}_{k},\mathbf{S}_{k},o_{k}$ is the parameters for the shape of each Gaussian. In the first stage, we utilize $\mathbf{c}_{k}$ as appearance to optimize the shape parameters of the Gaussians. At the second stage, to achieve PBR materials estimation and physically based rendering, we additionally introduce parameters albedo $\mathbf{a}_{k}\in\mathbb{R}^{3}$, roughness $r_{k}\in\mathbb{R}$ and metallics $m_{k}\in\mathbb{R}$ to model the decoupled appearance on each 2D Gaussian. We utilize $N_{\text{init}}=6980$ canonical SMPL template vertices and their corresponding normal vectors to initialize the location and rotation of our canonical Gaussians. The initialization of scaling factors is calculated in the canonical space following [40].
@@
Articulation via LBS. We utilize SMPL [39] as our articulation template model. The SMPL model is defined by a function $M(\boldsymbol{\theta},\boldsymbol{\beta}):\mathbb{R}^{|\boldsymbol{\theta}|\times|\boldsymbol{\beta}|}\rightarrow\mathbb{R}^{3N_{\text{init}}}$, where the $\boldsymbol{\theta}\in\mathbb{R}^{3\times K}$ ($K$ denotes the number of body joints) denotes the body pose parameters, $\boldsymbol{\beta}\in\mathbb{R}^{10}$ denotes the coefficients for body shape.
@@
With the template model, the Gaussian points can be deformed to the articulated space via linear blend skinning (LBS). For each canonical space Gaussian point $\mathbf{p}_{k}$, we first calculate the articulation transformation $\mathbf{A}_{k}=[\mathbf{A}_{k}^{rot},\mathbf{A}_{k}^{t}]$ based on the pose $\boldsymbol{\theta}$ where $\mathbf{A}_{k}^{rot}$ denotes the rotation matrix and $\mathbf{A}_{k}^{t}$ denotes the translation vector for the articulated transformation. The $\mathbf{A}_{k}$ for each Gaussian can be calculated via the following equation:
$\mathbf{A}_{k}=\sum_{i=1}^{N_{b}}\widehat{\mathbf{W}}_{i}(\mathbf{p}_{k})\mathbf{B}_{i}$ (5)
@@
$\mathcal{B}(\boldsymbol{\theta})=[\mathbf{B}_{1},\mathbf{B}_{2},\dots,\mathbf{B}_{N_{b}}]$ ($\mathbf{B}_{i}\in SE(3)$) denotes the rigid transformation that moves each canonical joint to the articulated space. $N_{b}$ is the number of joints in a template model. $\widehat{\mathbf{W}}(\mathbf{p}_{k})$ denotes the blending weight for the bones $\mathcal{B}(\boldsymbol{\theta})$. We model with $\widehat{\mathbf{W}}(\mathbf{p}_{k})=\mathbf{W}(\mathbf{p}_{k})+\Delta\mathbf{W}(\mathbf{p}_{k})$. $\mathbf{W}(\mathbf{p}_{k})\in R^{N_{b}}$ is the predefined skinning weight queried in canonical space for each Gaussian, and the $\Delta\mathbf{W}(\mathbf{p}_{k})$ is the learnable blending weight offset for each Gaussian. We follow [40] and add this term to the predefined blending $\mathbf{W}$ to model the actual instance deformation. We follow Instant Avatar [35] to store the predefined skinning weight of canonical space in a dense 3D volume. Then, we can deform the 2D Gaussians to the articulated space via:
$\mathbf{p}_{k}^{*}=\mathbf{A}_{k}^{rot}\mathbf{p}_{k}+\mathbf{A}_{k}^{t},\quad\mathbf{R}_{k}^{*}=\mathbf{A}_{k}^{rot}\mathbf{R}_{k}.$ (6)
@@
where $\mathbf{p}_{k}^{*}$ and $\mathbf{R}_{k}^{*}$ denotes the location and rotation of 2D Gaussians in articulated space.
@@
To approximate the clothing motion that can be captured from the monocular video, we follow [40] and introduce $N_{\text{lb}}=4$ latent Bones ~$\widetilde{\mathcal{B}}(\boldsymbol{\theta})$ parameterized with an MLP conditioned on human pose $\boldsymbol{\theta}$. We also parameterized the blending weight $\widetilde{\mathbf{W}}_{i}$ for each latent bones ~$\widetilde{\mathcal{B}}_{k}(\boldsymbol{\theta})$. We therefore extend the equation 5 to the overall articulation transformation:
$\mathbf{A}_{k}=\sum_{i=1}^{N_{b}}\widehat{\mathbf{W}}_{i}(\mathbf{p}_{k})\mathbf{B}_{i}+\sum_{i=1}^{N_{lb}}\widetilde{\mathbf{W}}_{i}(\mathbf{p}_{k})\widetilde{\mathbf{B}}_{i}$ (7)
3.3 Physically based rendering with image-based Lighting
@@
Physically Based Inverse Rendering. Unlike previous physically-based cloth avatars [29, 30], our lighting calculation does not need the time-consuming Monte Carlo Sampling, the integration of ingoing radiance in the hemisphere. We instead utilize the image-based lighting and split-sum approximation [101] to tackle the intractable integral.
The rendering equation involves the calculation of the outgoing light from a surfel at observation space $\mathbf{p}_{k}^{*}$  [^1] We omit $\mathbf{p}_{k}$ for simplified representation. All the $\mathbf{L}_{o},\mathbf{L}_{i},\mathbf{n},\mathbf{a},m$ in the following equation denotes the corresponding parameters on each surfel $\mathbf{p}^{*}_{k}$ based on the BRDF materials, which can be written as:
[…]
@@
where $\mathbf{w}_{o}$ denotes the view direction of outgoing radiance $\mathbf{L}_{o}(\mathbf{w}_{o})$, $\mathbf{w}_{i}$ denotes the direction of the ingoing radiance $\mathbf{L}_{i}(\mathbf{w}_{i})$ sampled from the hemisphere $\Omega$, $\mathbf{n}$ denotes the normal orientation of the surfel. The BRDFs property $f$ can be divided into a specular term (the 1st term in equation. 9) and a diffuse term (the 2nd term in equation. 9) following the Disney-BRDF model [102]:
$f(\mathbf{w}_{o},\mathbf{w}_{i})=\frac{DFG}{4(\mathbf{w}_{i}\cdot\mathbf{n})(\mathbf{w}_{o}\cdot\mathbf{n})}+(1-m)\frac{\mathbf{a}}{\pi}$ (9)
@@
where $D$ denotes the GGX [103] normal distribution function (NDF), $F$ denotes the Fresnel term, and $G$ denotes the geometric attenuation, $\mathbf{a}$ is the albedo for each Gaussian, $m$ denotes the metallics for each Gaussian. We can therefore separate the whole $\mathbf{L}_{o}(\mathbf{w}_{o})$ into two separate parts (specular light $\mathbf{L}_{s}$ in equation. 10 and diffuse light $\mathbf{L}_{d}$ in equation. 11). We utilize the split-sum approximation [101] to approximate the two terms separately. By approximation, we can separate the integral of the light and BRDF product into two separate integrals, which we can pre-integrate separately for rapid querying.
@@
[…]
The first term denotes the integral of the specular BRDFs. It only depends on the $\mathbf{w}_{i}\cdot\mathbf{n}$ and the roughness $r$ and is independent of the illumination. The second term denotes the integral of ingoing radiance with Specular NDF $D$, and it depends on the reflection direction $\mathbf{r}=2(\mathbf{w}_{i}\cdot\mathbf{n})\mathbf{n}-\mathbf{w}_{i}$ and roughness $r$. It can also be pre-integrated and queried from a filtered cubemap.
@@
[…]
@@
Occlusion Calculation. As there exists a lot of self-occlusion during the animation of clothed human avatars, to achieve more realistic physically based rendering, we should consider the occlusion. However, calculating the occlusion with ray-tracing [51] for each Gaussian is time-consuming and hinders rendering speed. To achieve occlusion calculation while also maintaining the fast rendering nature of Gaussian Splatting, we proposed to approximate the occlusion as the ambient occlusion from SMPL [39] mesh $M(\boldsymbol{\theta},\boldsymbol{\beta})=(\mathbf{V},\mathbf{F})$. Baking the ambient occlusion of the SMPL mesh can reduce the time for occlusion calculation. Specifically, we first query the nearest surface points $\mathbf{v}_{k}^{*}$ from mesh $M(\boldsymbol{\theta},\boldsymbol{\beta})$ for each 2D Gaussian $\mathbf{p}_{k}^{*}$ at articulated space. We then utilize the surface points $\mathbf{v}_{k}^{*}$ and its corresponding surface normal $\mathbf{n}_{k}^{*}$ from mesh to bake the ambient occlusion for the $k$-th Gaussian:
[…]
@@
$\mathbf{c}_{k}^{\text{PBR}}=\mathbf{L}_{o}(\mathbf{w}_{o})=\mathbf{L}_{d}+(1-AO_{k})\mathbf{L}_{s}$ (13)
here the $\mathbf{w}_{o}$ denotes the outgoing direction, corresponding to the view direction for each 2D gaussian in world space.
Environment Map Representation. We utilize a $6\times 64\times 64$ trainable parameters as the environment map for Image-based lighting. We query the light at different mip-level. The base level signifies the pre-integrated lighting for the minimum supported roughness value. Subsequent smaller mip-levels are derived from this base level using pre-filtering approach [101].
@@
Physical based Rendering with splatting. We first project the 2D Gaussians into image space with the camera parameters. We can then calculate the PBR color $\mathbf{c}_{k}^{PBR}$ for each 2D Gaussian and apply the splatting as
$\mathbf{c}^{\text{PBR}}(x)=\sum_{i=1}\mathbf{c}_{i}^{\text{PBR}}o_{i}\mathcal{G}_{i}(\mathbf{u}(x))\prod_{j=1}^{i-1}(1-o_{j}\mathcal{G}_{j}(\mathbf{u}(x)))$ (14)
we can also get the corresponding albedo, roughness, metallic, and visibility maps by following the same splatting process of 2DGS.

### v3

@@ 3 Our Method
Figure 2: Overview of our pipeline. We define the radiance ($c_{k}$) and materials ($a_{k},r_{k},m_{k}$) at the canonical space as canonical PBR-aware 2DGS and deform them to the world space via Linear Blend Skinning (LBS). We first optimize Image reconstruction loss to get the initial shape of the clothed avatar. Based on the rough shape, we optimize the Gaussian attributes with Physically-Based Rendering. To model the shadow effect and decouple the materials from lighting, we propose to approximate the calculation of occlusion with the template mesh. Additionally, we apply regularization, i.e., smooth loss, white loss, and normal consistency loss to get a plausible solution for the PBR materials.
In this section, we introduce our SGIA framework (shown in Fig.2) for fast and relightable dynamic clothed human reconstruction from monocular videos. Our method builds upon surfel-based representation[33] as described in Sec. 3.1. We begin by defining our proposed human representation in Sec. 3.2. Next, we introduce our occlusion calculation strategy, which enables efficient disentanglement of light and material properties, as detailed in Sec. 3.3. To ensure effective optimization in our framework, we then present a novel training strategy in Sec. 3.4. Finally, we describe the animation and relighting process in Sec. 3.5.
@@ 3.1 A revisit of 2DGS
Unlike 3DGS [32], 2DGS [33] adopts flat 2D Gaussians embedded in 3D space for scene representation. The primitive distributes densities within a planar disk (also known as surfel), which makes the primitive better align with surfaces. This feature can improve the quality of our geometry reconstruction and achieve better material estimation. In 2DGS [33], each primitive is characterized by a central point $\mathbf{p}_{k}\in\mathbb{R}^{3}$ , two principal tangential vectors $\mathbf{t}_{u},\mathbf{t}_{v}\in\mathbb{R}^{3}$ and two scaling factors $s_{u},s_{v}\in\mathbb{R}$. The scaling factors correspond to the variances of the 2D Gaussian.
@@
$\mathbf{P}(u,v)=\mathbf{p}_{k}+s_{u}\mathbf{t}_{u}u+s_{v}\mathbf{t}_{v}v=[\mathbf{R}_{k}\mathbf{S}_{k},\mathbf{p}_{k}](u,v,1,1)^{T}\text{,}$ (1)
where $\mathbf{R}_{k}=[\mathbf{t}_{u},\mathbf{t}_{v},\mathbf{t}_{u}\times\mathbf{t}_{v}]\in SO(3)$ defines the orientation of each surfel, $\mathbf{t}_{u}\times\mathbf{t}_{v}$ represents the surfel’s normal vector, and $\mathbf{S}_{k}=\text{diag}(s_{u},s_{v},0)\in\mathbb{R}^{3\times 3}$ is a diagonal matrix encoding the scale.
To render a pixel $x$, 2DGS accumulates the color by first calculating the intersection point between each ray and each tangent plane $\mathbf{P}$, and then accumulating the texture via volumetric alpha blending. The efficient ray-splat intersection calculation process involves finding the intersection between a ray and a local plane in the world space and the algorithm for this is detailed in the original paper [33].
For each intersection point $x$, we can transform it into the local coordinate defined by the tangent plane $\mathbf{P}$ as $\mathbf{u}=(u,v)$, and calculate its 2D Gaussian value using the following equation:
$\mathcal{G}(\mathbf{u})=\exp(-\frac{u^{2}+v^{2}}{2})\text{,}$ (2)
Then, the 2D Gaussians are sorted based on the depth of their centers and organized into tiles.
Finally, the color $\mathbf{c}(x)$ at the pixel $x$ can be calculated using the following equation:
$\mathbf{c}(x)=\sum_{i=1}^{n}\mathbf{c}_{i}o_{i}\mathcal{G}i(\mathbf{u}(x))\prod_{j=1}^{i-1}(1-o_{j}\mathcal{G}_{j}(\mathbf{u}(x)))\text{,}$ (3)
where $o_{i}$ denotes the opacity and $\mathbf{c}_{i}$ denotes the view-dependent appearance associated with the $i$-th 2D Gaussian.
@@ 3.2 Clothed humans avatars as dynamic surfels animated by template models
We build upon recent advancements in modeling humans as animatable Gaussians [39, 105, 43, 40] Specifically, we define the PBR-aware 2DGS in a canonical space , and transform them into world space using LBS. The Gaussians are animated by a human template model [38] to achieve avatar animation. We choose a surfel-based primitive [33] as our representation, as it provides improved geometry accuracy. Furthermore, we define additional parameters for intrinsic properties and lighting on each 2D Gaussian, enabling physically based rendering capabilities.
@@
$\mathbf{G}=\{\mathbf{p}_{k},\mathbf{R}_{k},\mathbf{S}_{k},\mathbf{c}_{k},o_{k},\mathbf{a}_{k},r_{k},m_{k}\}_{k=1}^{N_{\text{gs}}}\text{,}$ (4)
@@
Here, $\mathbf{p}_{k},\mathbf{R}_{k},\mathbf{S}_{k},o_{k}$ are the parameters that defined the shape of each Gaussian and $N_{\text{gs}}$ is the number of the gaussians. In the first stage, we optimize the shape parameters of the Gaussians using the appearance information encoded in $\mathbf{c}_{k}$. Then, to achieve physically-based materials estimation and rendering, we introduce additional parameters: albedo $\mathbf{a}_{k}\in\mathbb{R}^{3}$, roughness $r_{k}\in\mathbb{R}$ and metallics $m_{k}\in\mathbb{R}$. These parameters model the decoupled appearance properties on each 2D Gaussian. We initialize the location and rotation of our canonical Gaussians using $N_{\text{init}}=6980$ vertices and their corresponding normal vectors from the SMPL template model [38], The scaling factors are calculated in the canonical space following the approach proposed in [39].
@@
Articulation via LBS. We utilize the SMPL model [38] as our articulation template. The SMPL model is defined by a function $M(\boldsymbol{\theta},\boldsymbol{\beta}):\mathbb{R}^{|\boldsymbol{\theta}|\times|\boldsymbol{\beta}|}\rightarrow\mathbb{R}^{3N_{\text{init}}}$, where the $\boldsymbol{\theta}\in\mathbb{R}^{3\times K}$ denotes the body pose parameters. $\boldsymbol{\beta}\in\mathbb{R}^{10}$ denotes the coefficients for the body shape.
@@
With the SMPL template model, we can deform the Gaussian primitives to the articulated space using LBS. For each Gaussian point $\mathbf{p}_{k}$ in the canonical space, we first calculate the articulation transformation $\mathbf{A}_{k}=[\mathbf{A}_{k}^{rot},\mathbf{A}_{k}^{t}]$ , where $\mathbf{A}_{k}^{rot}$ denotes the rotation matrix and $\mathbf{A}_{k}^{t}$ denotes the translation vector. These articulation parameters are derived from the pose parameters $\boldsymbol{\theta}$ of the SMPL model. The transformation $\mathbf{A}_{k}$ for each Gaussian can be calculated using the following equation:
$\mathbf{A}_{k}=\sum_{i=1}^{N_{b}}\widehat{\mathbf{W}}_{i}(\mathbf{p}_{k})\mathbf{B}_{i}\text{,}$ (5)
@@
$\mathcal{B}(\boldsymbol{\theta})=[\mathbf{B}_{1},\mathbf{B}_{2},\dots,\mathbf{B}_{N_{b}}]$ ($\mathbf{B}_{i}\in SE(3)$) denotes the rigid transformation that moves each canonical joint to the articulated space. $N_{b}$ is the number of joints in a template model. $\widehat{\mathbf{W}}(\mathbf{p}_{k})$ denotes the blending weight for the bones $\mathcal{B}(\boldsymbol{\theta})$. We model with $\widehat{\mathbf{W}}(\mathbf{p}_{k})=\mathbf{W}(\mathbf{p}_{k})+\Delta\mathbf{W}(\mathbf{p}_{k})$. $\mathbf{W}(\mathbf{p}_{k})\in R^{N_{b}}$ is the predefined skinning weight queried in canonical space for each Gaussian, and the $\Delta\mathbf{W}(\mathbf{p}_{k})$ is the learnable blending weight offset for each Gaussian. Building upon the work of [39], we add an additional term to the predefined blending weights $\mathbf{W}$ to model the actual instance deformation. Then, we can deform the 2D Gaussians to the articulated space via:
$\mathbf{p}_{k}^{*}=\mathbf{A}_{k}^{rot}\mathbf{p}_{k}+\mathbf{A}_{k}^{t},\quad\mathbf{R}_{k}^{*}=\mathbf{A}_{k}^{rot}\mathbf{R}_{k}\text{,}$ (6)
@@
where $\mathbf{p}_{k}^{*}$ and $\mathbf{R}_{k}^{*}$ denotes the location and rotation of a 2D Gaussian in articulated space.
@@
To approximate the clothing motion captured from the monocular video, we follow the approach proposed in [39] and introduce $N_{\text{lb}}=4$ latent Bones $\widetilde{\mathcal{B}}(\boldsymbol{\theta})$ parameterized using a MLP conditioned on the human pose $\boldsymbol{\theta}$. Additionally, we parameterize the blending weight $\widetilde{\mathbf{W}}_{i}$ for each latent bones $\widetilde{\mathcal{B}}_{k}(\boldsymbol{\theta})$. This allows us to extend the articulation transformation equation. 5 to the overall articulation:
$\mathbf{A}_{k}=\sum_{i=1}^{N_{b}}\widehat{\mathbf{W}}_{i}(\mathbf{p}_{k})\mathbf{B}_{i}+\sum_{i=1}^{N_{lb}}\widetilde{\mathbf{W}}_{i}(\mathbf{p}_{k})\widetilde{\mathbf{B}}_{i}\text{,}$ (7)
3.3 Physically-Based Rendering with image-based Lighting
@@
Physically-Based Inverse Rendering. Unlike previous physically-based cloth avatars [29, 30], our lighting calculation does not need the time-consuming Monte Carlo Sampling, the integration of ingoing radiance in the hemisphere. We instead utilize the image-based lighting and split-sum approximation [106] to tackle intractable integral.
The rendering equation calculates the outgoing light from a surfel at the observation space position $\mathbf{p}_{k}^{*}$  [^1] For simplicity, we will omit the subscript $k$ in the following equations, but note that all the terms - $\mathbf{L}_{o}$, $\mathbf{L}_{i}$, $\mathbf{n}$, $\mathbf{a}$, and $m$ correspond to the parameters associated with the specific surfel $\mathbf{p}^{*}_{k}$. The rendering equation can be written as:
$\mathbf{L}_{o}(\mathbf{w}_{o})=\int_{\Omega}f(\mathbf{w}_{o},\mathbf{w}_{i})\mathbf{L}_{i}(\mathbf{w}_{i})(\mathbf{w}_{i}\cdot\mathbf{n})d\mathbf{w}_{i}\text{,}$ (8)
@@
where $\mathbf{w}_{o}$ denotes the view direction of outgoing radiance $\mathbf{L}_{o}(\mathbf{w}_{o})$, $\mathbf{w}_{i}$ denotes the direction of the ingoing radiance $\mathbf{L}_{i}(\mathbf{w}_{i})$ sampled from the hemisphere $\Omega$, and $\mathbf{n}$ denotes the normal orientation of the surfel. The BRDFs property $f$ can be divided into a specular term (the 1st term in equation. 9) and a diffuse term (the 2nd term in equation. 9) following the Disney-BRDF model [107]:
$f(\mathbf{w}_{o},\mathbf{w}_{i})=\frac{DFG}{4(\mathbf{w}_{i}\cdot\mathbf{n})(\mathbf{w}_{o}\cdot\mathbf{n})}+(1-m)\frac{\mathbf{a}}{\pi}\text{,}$ (9)
@@
where $D$ denotes the GGX [108] normal distribution function (NDF), $F$ denotes the Fresnel term [108], and $G$ denotes the geometric attenuation, $\mathbf{a}$ is the albedo for each Gaussian, $m$ denotes the metallics for each Gaussian. We further optimize the rendering equation by separating the outgoing radiance $\mathbf{L}_{o}(\mathbf{w}_{o})$ into two distinct parts (specular light $\mathbf{L}_{s}$ in equation. 10 and diffuse light $\mathbf{L}_{d}$ in equation. 11). We utilize the split-sum approximation proposed in  [106]. This allows us to separate the integral of the light and BRDF product into two separate integrals, which can then be pre-integrated for efficient querying during rendering.
@@
$\mathbf{L}_{s}\approx\int_{\Omega}\frac{DFG}{4(\mathbf{w}_{i}\cdot\mathbf{n})(\mathbf{w}_{o}\cdot\mathbf{n})}d\mathbf{w}_{i}\int_{\Omega}\mathbf{L}_{i}(\mathbf{w}_{i})D(\mathbf{w}_{i},\mathbf{w}_{o})(\mathbf{w}_{i}\cdot\mathbf{n})d\mathbf{w}_{i}\text{.}$ (10)
The first term in the specular light equation represents the integral of the specular BRDF, which depends only on the dot product $\mathbf{w}_{i}\cdot\mathbf{n}$ and the roughness $r$, independent of the illumination. The second term corresponds to the integral of the incoming radiance and the specular normal distribution function (NDF) $D$. It depends on the reflection direction $\mathbf{r}=2(\mathbf{w}_{i}\cdot\mathbf{n})\mathbf{n}-\mathbf{w}_{i}$ and the roughess $r$, which can be pre-integrated and queried from a filtered cubemap.
@@
$\mathbf{L}_{d}\approx\mathbf{a}(1-m)\int_{\Omega}L(\mathbf{w}_{i})\frac{\mathbf{w}_{i}\cdot\mathbf{n}}{\pi}d\mathbf{w}_{i}\text{.}$ (11)
@@
Occlusion Calculation. As there exists a lot of self-occlusion during the animation of clothed human avatars, to achieve more realistic physically based rendering, we should consider the occlusion. However, calculating the occlusion with ray-tracing [50] for each Gaussian is time-consuming and hinders rendering speed. To achieve occlusion calculation while also maintaining the fast rendering nature of Gaussian Splatting, we proposed to approximate the occlusion as the ambient occlusion from a SMPL [38] mesh $M(\boldsymbol{\theta},\boldsymbol{\beta})=(\mathbf{V},\mathbf{F})$. Baking the ambient occlusion of the SMPL mesh can reduce the time for occlusion calculation. Specifically, we first query the nearest surface points $\mathbf{v}_{k}^{*}$ from the mesh $M(\boldsymbol{\theta},\boldsymbol{\beta})$ for each 2D Gaussian $\mathbf{p}_{k}^{*}$ at articulated space. We then utilize the surface points $\mathbf{v}_{k}^{*}$ and its corresponding surface normal $\mathbf{n}_{k}^{*}$ from the mesh to bake the ambient occlusion for the $k$-th Gaussian:
$AO_{k}=\frac{1}{\pi}\int V_{\mathbf{v}_{k}^{*},\mathbf{w}}(\mathbf{n}_{k}^{*}\cdot\mathbf{w})d\mathbf{w}\text{,}$ (12)
@@
$\mathbf{c}_{k}^{\text{PBR}}=\mathbf{L}_{o}(\mathbf{w}_{o})=\mathbf{L}_{d}+(1-AO_{k})\mathbf{L}_{s}\text{,}$ (13)
where the $\mathbf{w}_{o}$ denotes the outgoing direction, corresponding to the view direction for each 2D gaussian in world space.
Environment Map Representation. We use the $6\times 64\times 64$ tensor of trainable parameters to represent the image-based lighting. We query the lighting information at different mipmap levels. The base level corresponds to the pre-integrated lighting for the minimum supported roughness value. The subsequent smaller mipmap levels are derived from this base level using a pre-filtering approach, as described in  [106].
@@
Physical-Based Rendering with splatting. To render the final color, we first calculate the ray-splat intersections for each pixel and Gaussian primitive. We then apply the following accumulation equation:
$\mathbf{c}^{\text{PBR}}(x)=\sum_{i=1}\mathbf{c}_{i}^{\text{PBR}}o_{i}\mathcal{G}_{i}(\mathbf{u}(x))\prod_{j=1}^{i-1}(1-o_{j}\mathcal{G}_{j}(\mathbf{u}(x)))\text{.}$ (14)
we can also get the corresponding albedo, roughness, metallic, and visibility maps following the same splatting process of 2DGS.